Western Air sold twenty-seven airline tickets to Cape Town on a certain day. Each Economy class ticket (E) was sold at N$3200 and the First-class tickets (F) were sold at N$4100 each. The total amount received for the twenty-seven tickets was N$92700.
4.5.1 Formulate the above information into two linear equations [2]
4.5.2 Solve these equations simultaneously using the substitution method and calculate the number of first-class tickets sold. [6]
1
Expert's answer
2020-05-19T18:26:04-0400

4.5.1 Let "x" be the number of Economy class tickets and "y" be the number of the First-class tickets. The total number will be

"x+y=27" .

The total amount of money will be

"3200\cdot x + 4100\cdot y = 92700."

Therefore, the linear system is

"\begin{cases}
x+y=27 \\
3200\cdot x + 4100\cdot y = 92700.
\end{cases}"

4.5.2 Let us divide the second equation by 100, for convenience, so we get

"\begin{cases}
x+y=27 \\
32\cdot x + 41\cdot y = 927.
\end{cases}"

Next, we express "x" from the first equation and substitute it into the second:

"\begin{cases}
x=27-y \\
32\cdot(27-y) + 41\cdot y = 927.
\end{cases}"

We should solve the second equation

"\begin{cases}
x=27-y \\
32\cdot27 + 9\cdot y = 927.
\end{cases}"

"\begin{cases}
x = 27-y,\\
y=7.
\end{cases}"

"\begin{cases}
x = 20,\\
y=7.
\end{cases}"

Therefore, the number of First-class tickets is 7.